Two charged objects in a laboratory are brought to 0.22 meters away from each other. If the force between the spheres is 4,550 N and the charge on one object is 9.9 × 10-5 C, what is the charge on the other object?
1
Expert's answer
2020-12-04T07:33:17-0500

The force acting between two charged objects (Coulomb force) is

"\displaystyle F = \frac{k q_1 q_2}{r^2}"

"\displaystyle q_2 = \frac{F r^2}{kq_1} = \frac{4550 \cdot0.22^2}{9 \cdot 10^9 \cdot 9.9 \cdot 10^{-5}} = 2.47 \cdot 10^{-4} \; C"

Answer: "2.47 \cdot 10^{-4} \; C"